Find the x-intercepts. y=x^2-4x+4.
Answer by scott8148(6628)

the x-intercepts occur when the graph intercepts the x-axis, at this point y=0

set y=0 and solve for x ... 0=x^2-4x+4 ... factoring gives 0=(x-2)(x-2) ... so x=2

the fact that there is only one intercept value means that the graph is tangent to the x-axis